OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

@PinkSapphire @Jagking

OpenStudy (anonymous):

75

OpenStudy (anonymous):

i think.

OpenStudy (pinksapphire):

C=81.09

OpenStudy (anonymous):

wht?!

OpenStudy (anonymous):

Oh yeah. @PinkSapphire is right.

OpenStudy (anonymous):

all i did was add up the discounts and subtracted tht from the orignal cost.

OpenStudy (anonymous):

x=the original price which is 100

OpenStudy (pinksapphire):

Ha! @Jagking

OpenStudy (anonymous):

then you substitute 100 for x, right?

OpenStudy (pinksapphire):

Just plug in 100 in for x in the equation.

OpenStudy (anonymous):

k. watch this. whts next kayla?

OpenStudy (pinksapphire):

Yes.

OpenStudy (anonymous):

Okay...I got it know. lol I didnt realize how easy that one was till I retried it.

OpenStudy (anonymous):

Can yall show me how to do this one? Evan is making a table that will be created in the shape of the figure below. The table top is a triangle attached to a rectangle. To purchase the right amount of paint, he needs to know the area of the table top. He can only spend $10 on paint, which is enough to cover 150 ft2 of surface area. What is the maximum length of the base of the rectangle he can build? An image of a compound shape made up of a rectangle and a triangle is shown. The length of the rectangle is labeled x and the width is labeled 6 feet. The base of the triangle is labeled 4 feet and the height is labeled 6 feet. x = __________________ ft

OpenStudy (anonymous):

This is the picture it shows

OpenStudy (anonymous):

dang it. i wasn't ready. i had 2 go somewhere. oh. i'm not good at word problems. i'll let pinky get this.

OpenStudy (pinksapphire):

Thanks:)

OpenStudy (anonymous):

>:(

OpenStudy (anonymous):

but u better hurry pinky. i'm working on it.

OpenStudy (anonymous):

@Jagking Can you help me with a different on then?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

Let the integers be x + (x+1) + (x + 2) x + (x+1) + (x + 2) = 267 3x + 3 = 267 3x = 264 x = 88 88+89+90=267 Largest integer=90 Another method 267/3 = 89 It's the average,so it has to be the middle one. Thus 89, then I took 89 - 1 = 88 and 89 + 1 = 90 and just put them in order. 88+89+90=267 Largest integer=90

OpenStudy (pinksapphire):

The area of the triagle is 12...I'm not done yet.

OpenStudy (anonymous):

anymore kalya?

OpenStudy (pinksapphire):

x=23

OpenStudy (anonymous):

wow. i was just getting ready 2 go 2 sleep..

OpenStudy (anonymous):

pinky.

OpenStudy (pinksapphire):

2 bad.

OpenStudy (anonymous):

u need more help @kaylamarie101

OpenStudy (anonymous):

Only one more

OpenStudy (anonymous):

the rest I already completed

OpenStudy (anonymous):

As the number of pages in a photo book increases, the price of the book also increases. There is an additional shipping charge of 15%. The price of a book can be modeled by the equation below where, P = the price of the book, 20 is the printing charge, 0.5 is the charge per page, and x = the number of pages. P = (20 + 0.5x) + 0.15(20 + 0.5x) Jennifer wants to purchase a book but only has $62.10 to spend. What is the maximum number of pages she can have in her book? x = ______________________ pages

OpenStudy (anonymous):

x=68

OpenStudy (anonymous):

is pinky giving up?

OpenStudy (anonymous):

Thank You...How did you get that answer? @Jagking And @PinkSapphire how did you get x=23 for before?

OpenStudy (anonymous):

i didn't do tht one. tht was pinky.

OpenStudy (anonymous):

you got x=68 for the last one, but how?

OpenStudy (anonymous):

oh. i made the equation equal 2: (20 + 0.5x) + 0.15(20 + 0.5x) and io sovled 4 x.

OpenStudy (anonymous):

Oh, okay. @PinkSapphire How did you get x=23

OpenStudy (anonymous):

And thanks again jag

OpenStudy (anonymous):

ur welcome. now pinky......

OpenStudy (pinksapphire):

12+6x=150 -12 -12 6x=138 6x/6=138/6 x=23

OpenStudy (anonymous):

Okay...Thanks @PinkSapphire